# Complexity of Abduction in Łukasiewicz Logic

### v1

@@ 4.3 Relevance and Necessity of Hypotheses
$\Gamma^{\mathsf{r}}$ =$=\{t\!\rightarrow\!\psi\mid\psi\!\in\!\Gamma\}\!\cup\!\{t^{\prime}\!\rightarrow\!\chi,t\!\rightarrow\!\neg t^{\prime},t\!\rightarrow\!t^{\prime\prime},t^{\prime}\!\rightarrow\!t^{\prime\prime}\}$
@@
$\mathcal{S}(\mathbb{P}^{\mathsf{r}})$ =$=\left\{\!\varrho\!\odot\!(t\!\geq\!\mathbf{1})\!\mid\!\varrho\!\in\!\mathcal{S}(\mathbb{P}^{{\textsf{\L}}})\right\}\!\cup$
@@ 5.1 Simple Clause Fragment
Let us now tackle the hardness. We provide a logspace reduction from (proper, entailment-minimal) solution recognition for classical Horn theories, which is ${\mathsf{P}}$-complete. Let $\mathbb{P}=\langle\Gamma,p,{\mathsf{H}}\rangle$ be a classical Horn abduction problem and $p\notin{\mathsf{H}}$, clauses in $\Gamma$ be finite sets of literals, and $\tau\in\mathcal{S}(\mathbb{P})$. I.e., $\Gamma\cup\{\tau\}$ is $\mathsf{CPL}$-satisfiable, and there is a unit resolution inference of $\{p\}$ from $\Gamma\cup\{\tau\}$ (i.e., in each application of the resolution rule, at least one premise is a unit clause). Let $\Gamma^{\oplus}$ be the result of replacing each clause $\{l_{1},\ldots,l_{k}\}$ in $\Gamma$ with $l_{1}\oplus\ldots\oplus l_{k}$; $\tau^{\lozenge{{\textsf{\L}}}}$ be the result of replacing $\wedge$ with $\odot$, positive literals $r$ in $\tau$ with $r\!\geq\!\mathbf{1}$, and negative literals $\neg q$ with $q\leq\mathbf{0}$; and ${\mathsf{H}}^{{\textsf{\L}}}$ be the result of replacing $r$ in ${\mathsf{H}}$ with $r\!\geq\mathbf{1}$, and $\neg q$ with $q\!\leq\!\mathbf{0}$. The size of $\mathbb{P}^{\oplus}=\langle\Gamma^{\oplus},p,{\mathsf{H}}^{{\textsf{\L}}}\rangle$ is linear in the size of $\mathbb{P}$, and $\Gamma^{\oplus}\cup\{\tau^{\lozenge{{\textsf{\L}}}}\}$ is Ł-satisfiable because $\Gamma\cup\{\tau\}$ is $\mathsf{CPL}$-satisfiable. Furthermore, unit resolution is Ł-sound: if $v(l_{1}\oplus\ldots\oplus l_{k})=1$ and $v(l_{1})=0$, then $v(l_{2}\oplus\ldots\oplus l_{k})=1$. Hence, reusing the inference in Ł, we have $\Gamma^{\oplus}\cup\{\tau^{\lozenge{{\textsf{\L}}}}\}\models_{{\textsf{\L}}}p$. Thus, $\tau^{\lozenge{{\textsf{\L}}}}\in\mathcal{S}(\mathbb{P}^{\oplus})$. Conversely, let $\tau$ solve $\mathbb{P}^{\oplus}$, i.e., $\Gamma^{\oplus}\cup\{\tau^{\lozenge{{\textsf{\L}}}}\}\models^{\mathsf{cons}}_{{\textsf{\L}}}p$. Let $\tau_{\mathsf{cl}}$ be obtained from $\tau$ by replacing $\odot$ with $\wedge$, $r\!\geq\!\mathbf{1}$ with $r$, and $q\!\leq\!\mathbf{0}$ with $\neg q$. Clearly, $\Gamma,\tau_{\mathsf{cl}}\models_{\mathsf{CPL}}p$. Assume for contradiction that $\Gamma,\tau_{\mathsf{cl}}\models_{\mathsf{CPL}}\bot$. Then there is a classical unit resolution inference of the empty clause from $\Gamma\cup\{\tau_{\mathsf{cl}}\}$. As it is sound in Ł, $\Gamma,\tau_{\mathsf{cl}}\models_{{\textsf{\L}}}\bot$. Contradiction.
@@ 5.2 Interval Clause Fragment
Results in Section 4 show that the complexity of Ł-abduction in the general case was not affected by whether we allow the use of interval literals in theories or observations (recall from Definition 4 that theories and observations are defined over $\mathcal{L}^{\mathbb{Q}}_{{\textsf{\L}}}$). Interval literals can also reduce the size and facilitate the understanding (for a human) of an abduction problem while preserving its solutions. E.g., the theory $\Gamma_{\mathsf{lift}}$ of the problem $\mathbb{P}_{\mathsf{lift}}$ from Examples 1 and 2 can be reformulated as follows: $\Gamma^{\mathbb{Q}}_{\mathsf{lift}}=\{(c\geq\mathbf{\tfrac{1}{4}})\leftrightarrow g,(c\leq\mathbf{\tfrac{2}{3}})\leftrightarrow b\}$. Observe that $(c\!\geq\!\mathbf{\tfrac{1}{4}})\!\leftrightarrow\!g$ is weakly equivalent to $(c\!\oplus\!c\!\oplus\!c\!\oplus\!c)\!\leftrightarrow\!g$ and $(c\!\leq\!\mathbf{\tfrac{2}{3}})\!\leftrightarrow\!b$ to ($(\neg c\!\oplus\!\neg c\!\oplus\!\neg c)\!\leftrightarrow\!b$. So, for $\mathbb{P}^{\mathbb{Q}}_{\mathsf{lift}}=\langle\Gamma^{\mathbb{Q}}_{\mathsf{lift}},g\odot b,{\mathsf{H}}_{\mathsf{lift}}\rangle$, we have $\mathcal{S}(\mathbb{P}_{\mathsf{lift}})=\mathcal{S}(\mathbb{P}^{\mathbb{Q}}_{\mathsf{lift}})$ (and likewise for other types of solutions). Moreover, using interval literals in $\oplus$-clauses simplifies the presentation of relations between values. E.g., ‘if $v(p)\geq\tfrac{1}{2}$, then $v(q)\!\leq\!\tfrac{3}{4}$’ can be put as $(p\!\geq\!\mathbf{\tfrac{1}{2}})\!\rightarrow\!(q\!\leq\!\mathbf{\tfrac{3}{4}})$.
@@
Membership follows from Theorem 1. Hardness can be shown via a reduction from the classical solution recognition. Namely, let $\mathbb{P}=\langle\Gamma,p,{\mathsf{H}}\rangle$ a classical abduction problem with $\Gamma$ being a set of $\vee$-clauses, we replace all positive literals $r$ occurring in $\mathbb{P}$ with $r\geq\mathbf{1}$ and all negative literals $\neg s$ with $s<\mathbf{1}$. Similarly, given a classical solution $\tau$, we produce an Ł-solution by a similar replacement of literals and by changing $\wedge$ to $\odot$. For the converse direction, given a solution for the Ł-abduction problem, a classical solution for $\mathbb{P}$ can be obtained by a reverse replacement. ∎
The $\Sigma^{\mathsf{P}}_{2}$-hardness of solution existence for Ł-abduction in the interval clause fragment can be obtained by a reduction from Theorem 12.
@@
As expected, determining the relevance of $\lambda\in{\mathsf{H}}$ is $\Sigma^{\mathsf{P}}_{2}$-hard. The next statement can be obtained by a reduction shown in (6). The only difference is that instead of variables $t$, $t^{\prime}$, $t^{\prime\prime}$, we take interval literals $t\geq\mathbf{1}$, $t^{\prime}\geq\mathbf{1}$, and $t^{\prime\prime}\geq\mathbf{1}$.
@@
On the other hand, it is possible to restrict the interval clause theories so that their complexity remains the same as that of the simple clause theories. Namely, we prohibit occurrences of pairs of literals $\lambda$ and $\lambda^{\prime}$ s.t. $\neg\lambda\oplus\neg\lambda^{\prime}$ is Ł-valid, e.g., $p\geq\mathbf{\nicefrac{{2}}{{3}}}$ and $p<\mathbf{\nicefrac{{3}}{{4}}}$.
@@
A set of interval clauses $\Gamma$ is cover-free (CF) if no pair of literals $\lambda$ and $\lambda^{\prime}$ s.t. Ł¬${{\textsf{\L}}}\models\neg\lambda\oplus\neg\lambda^{\prime}$ occurs in $\Gamma$.
@@ 6 Conclusion and Discussion
Finally, we recall that there are algorithms for solving fuzzy logic abduction problems with fuzzy sets or by specifying exact values of hypotheses. In our framework, however, a solution is an interval term corresponding to a set of intervals of permitted values of hypotheses. Thus, it would be instructive to devise algorithms for the generation of solutions in the form of interval terms. As satisfiability and validity in Ł are reducible to mixed-integer programming (MIP) (???), it would make sense to apply MIP solvers. Moreover, as the semantics of Ł can be given over rational numbers (?), one may try to apply rational MIP solvers such as the one by (?) as they produce exact solutions to mixed-integer problems.
@@ Acknowledgements
The $1^{\mathsf{st}}$ author was supported by…We would also like to thank the reviewers for their constructive suggestions and spotting a mistake in the first version of the paper.
@@ Appendix B Proofs of Section 4
$(p\geq\mathbf{c})^{\flat}_{\mathsf{H}}$ =$=\begin{cases}p\geq\mathbf{c}^{\downarrow}&\text{if }p\geq\mathbf{c}^{\downarrow}\in{\mathsf{H}}\\ p>\mathbf{c}^{\downarrow}&\text{otherwise}\end{cases}$
$(p>\mathbf{c})^{\flat}_{\mathsf{H}}$ =$=\begin{cases}p\geq\mathbf{c}&\text{if }p\geq\mathbf{c}\in{\mathsf{H}}\\ p\geq\mathbf{c}^{\downarrow}&\text{if }p\geq\mathbf{c}^{\downarrow}\in{\mathsf{H}}\\ p>\mathbf{c}^{\downarrow}&\text{otherwise}\end{cases}$
$(p\leq\mathbf{d})^{\flat}_{\mathsf{H}}$ =$=\begin{cases}p\leq\mathbf{d}^{\uparrow}&\text{if }p\geq\mathbf{d}^{\downarrow}\in{\mathsf{H}}\\ p>\mathbf{d}^{\uparrow}&\text{otherwise}\end{cases}$
$(p<\mathbf{d})^{\flat}_{\mathsf{H}}$ =$=\begin{cases}p\leq\mathbf{d}&\text{if }p\leq\mathbf{d}\in{\mathsf{H}}\\ p\leq\mathbf{d}^{\uparrow}&\text{if }p\leq\mathbf{d}^{\downarrow}\in{\mathsf{H}}\\ p<\mathbf{d}^{\uparrow}&\text{otherwise}\end{cases}$
@@
$\Gamma$ =$=\left\{p\vee\neg p\mid p\in\mathsf{Pr}(\chi\wedge q)\right\}\cup\{q\}$
${\mathsf{H}}$ =$=\{p\geq\mathbf{1}\mid p\in\mathsf{Pr}(\chi)\}\cup\{p\leq\mathbf{0}\mid p\in\mathsf{Pr}(\chi)\}$
@@
$v^{\mathsf{cl}}(r)$ =$=\begin{cases}1&\text{if }v(s\geq\mathbf{1})=1\text{ or }v(r\leq\mathbf{0})=0\\ 0&\text{if }v(r\leq\mathbf{0})=1\text{ or }v(r\geq\mathbf{1})=0\end{cases}$
@@
$\tau^{\flat}_{\lambda_{k}}$ =$=\begin{cases}\bigodot\limits_{{\begin{matrix}\lambda^{\prime}\in\tau\\ \lambda^{\prime}\neq\lambda\end{matrix}}}\lambda^{\prime}\odot{\lambda^{\flat}_{k}}_{\mathsf{H}}&\text{if }{\lambda^{\flat}_{k}}_{\mathsf{H}}\text{ is defined}\\ \bigodot\limits^{k-1}_{i=1}\lambda_{i}\odot\bigodot\limits^{n}_{j=k+1}\!\!\!\!\lambda_{j}&\text{otherwise}\end{cases}$
@@ Appendix C Proofs of Section 5
$\Gamma$ =$=\{\kappa_{i}\mid i\in\{1,\ldots,m\}\}\cup\{\tau_{j}\mid j\in\{1,\ldots,n\}\}$
@@
$\Gamma^{\mathbb{LI}}$ =$=\{\kappa^{\mathbb{LI}}_{i}\!\mid i\!\in\!\{1,\ldots,m\}\}\!\cup\!\bigcup\limits^{n}_{j=1}\{(\lambda^{j}_{i})^{\mathbb{LI}}\!\mid i\!\in\!\{1,\ldots,k_{j}\}\}$
@@
$\Gamma_{\mathbf{int}}$ =$=\left\{\bigoplus\limits_{q_{i}\in\kappa_{i}}(q_{i}\!\geq\!\mathbf{1})\oplus\bigoplus\limits_{\neg r_{i}\in\kappa_{i}}(r_{i}\!<\!\mathbf{1})\mid i\in\{1,\ldots,n\}\right\}$
${\mathsf{H}}_{\mathbf{int}}$ =$=\{s\geq\mathbf{1}\mid s\in{\mathsf{H}}\}\cup\{t<\mathbf{1}\mid\neg t\in{\mathsf{H}}\}$
@@
The proofs of Theorems 15 and 16 relies on the following lemma.
Lemma 1.
Let $\mathbb{P}=\langle\Gamma,\tau,{\mathsf{H}}\rangle$ be a classical abduction problem s.t. $\Gamma$ is a set of clauses containing only negative literals and $\tau$ a weak conjunction of literals. Then it is $\mathsf{NP}$-complete to decide whether $\mathcal{S}(\mathbb{P})=\varnothing$.
Proof.
The membership is evident because $\Gamma$ is a family of Horn clauses. For the hardness, consider the following reduction from $\mathsf{CPL}$-satisfiability that we adapt from (?, Theorem 5.3). Let $\Delta=\{\kappa_{1},\ldots,\kappa_{m}\}$ be a set of $\vee$-clauses and $\mathsf{Pr}[\Delta]=\{p_{1},\ldots,p_{n}\}$. Let further ${\mathsf{P}}^{\prime}=\{p^{\prime}_{1},\ldots,p^{\prime}_{n}\}$, ${\mathsf{Q}}=\{q_{1},\ldots,q_{m}\}$, and ${\mathsf{R}}=\{r_{1},\ldots,r_{n}\}$ be pair-wise disjoint sets of propositional variables that do not occur in $\Delta$. Define $\mathbb{P}^{\mathcal{H}}=\langle\Delta^{\mathcal{H}},\tau^{\mathcal{H}},{\mathsf{H}}^{\mathcal{H}}\rangle$
$\Delta^{\mathcal{H}}$ =$=\{\neg p_{i}\!\vee\!\neg p^{\prime}_{i}\mid 1\!\leq\!i\!\leq\!n\}\!\cup\!\{\neg p_{i}\!\vee\!\neg r_{i}\mid 1\!\leq\!i\!\leq\!n\}\cup$
$\quad~{}\{\neg p^{\prime}_{i}\!\vee\!\neg r_{i}\mid 1\!\leq\!i\!\leq\!n\}\cup$
$\quad~{}\bigcup\limits^{m}_{i=1}\left(\{\neg p_{j}\!\vee\!\neg q_{i}\mid p_{j}\!\in\!\kappa_{i}\}\!\cup\!\{\neg p^{\prime}_{j}\!\vee\!\neg q_{i}\mid\neg p_{j}\!\in\!\kappa_{i}\}\right)$
$\tau^{\mathcal{H}}$ $=\bigwedge\limits^{m}_{i=1}\neg q_{i}\wedge\bigwedge\limits^{n}_{j=1}\neg r_{j}$
${\mathsf{H}}^{\mathcal{H}}$ $=\mathsf{Pr}[\Delta]\cup{\mathsf{P}}^{\prime}$
One can show that $\Delta$ is $\mathsf{CPL}$-satisfiable iff $\mathcal{S}(\mathbb{P}^{\mathcal{H}})\neq\varnothing$. Indeed, if a classical valuation $v$ satisfies $\Delta$, then a solution is $\bigwedge_{v(p_{i})=1}p_{i}\wedge\bigwedge_{v(p_{i})=0}p^{\prime}_{i}$. Conversely, for a solution $\sigma$, a satisfying valuation is as follows: $v(p_{i})\!=\!1$ iff $p_{i}\!\in\!\sigma$. ∎
@@
For the membership, we construct a reduction to the classical Horn abduction for which solution recognition is known to be polynomial. Let $\Gamma=\{\kappa_{1},\ldots\kappa_{m}\}$ and, in particular $\kappa_{i}=\bigoplus_{j=1}^{n_{i}}(p_{j}\lozenge\mathbf{c}_{j})$ for each $i\in\{1,\ldots,m\}$. Using Definition 1 as well as (3) and (7), each clause $\kappa_{i}$ can be rewritten as follows: $\kappa^{\odot}_{i}=\bigodot_{j=1}^{n_{i}}(p_{j}\blacklozenge\mathbf{c}_{j})\rightarrow\bot$. Note that $\kappa_{i}\equiv_{{\textsf{\L}}}\kappa^{\odot}_{i}$.
Now set $\Gamma^{\odot}=\{\kappa^{\odot}_{i}\mid\kappa_{i}\in\Gamma\}$ and let $\mathbb{P}^{\odot}=\langle\Gamma^{\odot},\tau,{\mathsf{H}}\rangle$. It is clear that the sets of (resp., proper, Ł-minimal) solutions of $\mathbb{P}$ and $\mathbb{P}^{\odot}$ coincide.
We consider the following classical abduction problem: $\mathbb{P}_{\mathcal{H}}=\langle\Gamma_{\mathcal{H}},\tau_{\mathcal{H}},{\mathsf{H}}_{\mathcal{H}}\rangle$ defined as described below. Given an interval literal $\lambda$ over $p$, let $r_{\lambda}$ be a fresh variable and set:
$\Gamma_{\mathcal{H}}$ =$=\left\{\bigwedge\limits^{n_{i}}_{j=1}r_{p_{j}\blacklozenge\mathbf{c}_{j}}\rightarrow\bot\mid i\in\{1,\ldots,m\}\right\}\cup$
$\quad~{}~{}\left\{r_{\lambda}\!\wedge\!r_{\lambda^{\prime}}\!\rightarrow\!\bot\left|\begin{matrix}{{\textsf{\L}}}\!\models\!\lambda\odot\lambda^{\prime}\rightarrow\bot\\ \lambda\text{ and }\lambda^{\prime}\text{ occur in }\mathbb{P}^{\odot}\end{matrix}\right.\right\}\cup$
$\quad~{}~{}\left\{r_{\lambda}\rightarrow r_{\lambda^{\prime}}\left|\begin{matrix}{{\textsf{\L}}}\models\lambda\rightarrow\lambda^{\prime}\\ \lambda\text{ and }\lambda^{\prime}\text{ occur in }\mathbb{P}^{\odot}\end{matrix}\right.\right\}$
@@
${\mathsf{H}}_{\mathcal{H}}$ =$=\{r_{\lambda}\mid\lambda\in{\mathsf{H}}\}$ (9)
As one can see, $\Gamma^{\odot}_{\mathcal{H}}$ is polynomial in the size of $\Gamma^{\odot}$ (and hence, in the size of $\Gamma$) and (recall Proposition 3) it takes polynomial time to determine whether $\lambda\rightarrow\lambda^{\prime}$ is Ł-valid and $\lambda\odot\lambda^{\prime}$ is Ł-unsatisfieable. Furthermore, $\Gamma^{\odot}_{\mathcal{H}}$ is a set of Horn formulas and only contains variables of the form $r_{\lambda}$ with $\lambda$ being an interval term occurring in $\mathbb{P}^{\odot}$. We show that $\sigma$ is a (proper) solution to $\mathbb{P}^{\odot}$ iff $\sigma_{\mathcal{H}}=\bigwedge\limits_{\lambda\in\sigma}r_{\lambda}$ is a (proper) solution to $\mathbb{P}_{\mathcal{H}}$.
Assume that $\sigma\!\in\!\mathcal{S}(\mathbb{P}^{\odot})$. We show that $\Gamma_{\mathcal{H}},\sigma_{\mathcal{H}}\!\models^{\mathsf{cons}}_{\mathsf{CPL}}\!\tau_{\mathcal{H}}$. Let $v$ be an Ł-valuation witnessing that $\Gamma^{\odot},\sigma\not\models_{{\textsf{\L}}}\bot$. We construct the classical valuation $v_{\mathcal{H}}$ as follows:
$v_{\mathcal{H}}(r_{\lambda})$ =$=\begin{cases}1&\text{if }v(\lambda)=1\\ 0&\text{otherwise}\end{cases}$ (10)

### v3

@@ 4.3 Relevance and Necessity of Hypotheses
$\Gamma^{\mathsf{r}}$ $=\{t\!\rightarrow\!\psi\mid\psi\!\in\!\Gamma\}\!\cup\!\{t^{\prime}\!\rightarrow\!\chi,t\!\rightarrow\!\neg t^{\prime},t\!\rightarrow\!t^{\prime\prime},t^{\prime}\!\rightarrow\!t^{\prime\prime}\}$
@@
$\mathcal{S}(\mathbb{P}^{\mathsf{r}})$ $=\left\{\!\varrho\!\odot\!(t\!\geq\!\mathbf{1})\!\mid\!\varrho\!\in\!\mathcal{S}(\mathbb{P}^{{\textsf{\L }}})\right\}\!\cup$
@@ 5.1 Simple Clause Fragment
[…]
@@ 5.2 Interval Clause Fragment
Results in Section 4 show that the complexity of Ł-abduction in the general case was not affected by whether we allow the use of interval literals in theories or observations (recall from Definition 4 that theories and observations are defined over $\mathcal{L}^{\mathbb{Q}}_{{\textsf{\L }}}$). Interval literals can also reduce the size and facilitate the understanding (for a human) of an abduction problem while preserving its solutions. E.g., the theory $\Gamma_{\mathsf{lift}}$ of the problem $\mathbb{P}_{\mathsf{lift}}$ from Examples 1 and 2 can be reformulated as follows: $\Gamma^{\mathbb{Q}}_{\mathsf{lift}}=\{(c\geq\mathbf{\tfrac{1}{4}})\leftrightarrow g,(c\leq\mathbf{\tfrac{2}{3}})\leftrightarrow b\}$. Observe that $(c\!\geq\!\mathbf{\tfrac{1}{4}})\!\leftrightarrow\!g$ is weakly equivalent to $(c\!\oplus\!c\!\oplus\!c\!\oplus\!c)\!\leftrightarrow\!g$ and $(c\!\leq\!\mathbf{\tfrac{2}{3}})\!\leftrightarrow\!b$ to $(\neg c\!\oplus\!\neg c\!\oplus\!\neg c)\!\leftrightarrow\!b$. So, for $\mathbb{P}^{\mathbb{Q}}_{\mathsf{lift}}=\langle\Gamma^{\mathbb{Q}}_{\mathsf{lift}},g\odot b,{\mathsf{H}}_{\mathsf{lift}}\rangle$, we have $\mathcal{S}(\mathbb{P}_{\mathsf{lift}})=\mathcal{S}(\mathbb{P}^{\mathbb{Q}}_{\mathsf{lift}})$ (and likewise for other types of solutions). Moreover, using interval literals in $\oplus$-clauses simplifies the presentation of relations between values. E.g., ‘if $v(p)\geq\tfrac{1}{2}$, then $v(q)\!\leq\!\tfrac{3}{4}$’ can be put as $(p\!\geq\!\mathbf{\tfrac{1}{2}})\!\rightarrow\!(q\!\leq\!\mathbf{\tfrac{3}{4}})$.
@@
Membership follows from Theorem 1. Hardness can be shown via a reduction from solution recognition. in $\mathsf{CPL}$. Let $\mathbb{P}=\langle\Gamma,p,{\mathsf{H}}\rangle$ be a classical abduction problem with $\Gamma$ a set of $\vee$-clauses. We replace all positive literals $r$ occurring in $\mathbb{P}$ with $r\geq\mathbf{1}$ and all negative literals $\neg s$ with $s<\mathbf{1}$. Given $\tau\in\mathcal{S}(\mathbb{P})$, we produce an Ł-solution by a similar replacement of literals and changing $\wedge$ to $\odot$. Conversely, a classical solution for $\mathbb{P}$ can be obtained from a solution for the Ł-abduction problem by a reverse replacement. ∎
The $\Sigma^{\mathsf{P}}_{2}$-hardness of solution existence for Ł-abduction in the interval clause fragment can be obtained by a reduction used in Theorem 12.
@@
As expected, determining the relevance of $\lambda\in{\mathsf{H}}$ is $\Sigma^{\mathsf{P}}_{2}$-hard. The next statement can be obtained by a reduction shown in (6). The only difference is that instead of variables $t$, $^{\prime}, t′′t^{\prime\prime}$, we take interval literals $t\geq\mathbf{1}$, $t^{\prime}\geq\mathbf{1}$, and $t^{\prime\prime}\geq\mathbf{1}$.
@@
Still, we can restrict interval clause theories so that abduction becomes $\mathsf{NP}$-complete. Namely, we prohibit interval literals ‘covering’ the entire $[0,1]$ interval (e.g., $p\!\leq\!\mathbf{\tfrac{2}{3}}$ and $p\!\geq\!\mathbf{\tfrac{1}{4}}$) in the implicative representation of interval clauses.
@@
Let $\Gamma$ be a set of interval clauses represented as $\bigodot^{m}_{i=1}(p_{i}\lozenge\mathbf{c}_{i})\!\rightarrow\!(q\lozenge\mathbf{d})$ and $\bigodot^{m}_{i=1}(p_{i}\lozenge\mathbf{c}_{i})\!\rightarrow\!\bot$. $\Gamma$ is cover-free (CF) if no pair of interval literals $\lambda$ and $\lambda^{\prime}$ over the same variable occurs in $\Gamma$ s.t. ${{\textsf{\L }}}\models\lambda\oplus\lambda^{\prime}$.
@@ 6 Conclusion and Discussion
Finally, recall that there are algorithms for solving fuzzy logic abduction problems with fuzzy sets. In our framework, however, a solution corresponds to a set of intervals of permitted values of hypotheses. As Ł-entailment is reducible to finite-valued Łukasiewicz logics (?), we might hope that it is possible to compute the ‘minimally sufficient’ set of permitted values from the shape of the theory. Thus, it would be instructive to devise algorithms for the generation of solutions in the form of interval terms. As satisfiability and validity in Ł are reducible to mixed-integer programming (MIP) (??), it would make sense to apply MIP solvers. Moreover, since the semantics of Ł can be given over rational numbers (?), one may try to apply rational MIP solvers such as the one by (?) as they produce exact solutions to mixed-integer problems.
@@ Acknowledgements
The $1^{\mathsf{st}}$ author was supported by JSPS KAKENHI Grant Number JP25K03190 and JST CREST Grant Number JPMJCR22D3, Japan. We would also like to thank the reviewers for their constructive suggestions and spotting a mistake in the first version of the paper.
@@ Appendix B Proofs of Section 4
$(p\geq\mathbf{c})^{\flat}_{\mathsf{H}}$ $=\begin{cases}p\geq\mathbf{c}^{\downarrow}&\text{if }p\geq\mathbf{c}^{\downarrow}\in{\mathsf{H}}\\ p>\mathbf{c}^{\downarrow}&\text{otherwise}\end{cases}$
$(p>\mathbf{c})^{\flat}_{\mathsf{H}}$ $=\begin{cases}p\geq\mathbf{c}&\text{if }p\geq\mathbf{c}\in{\mathsf{H}}\\ p\geq\mathbf{c}^{\downarrow}&\text{if }p\geq\mathbf{c}^{\downarrow}\in{\mathsf{H}}\\ p>\mathbf{c}^{\downarrow}&\text{otherwise}\end{cases}$
$(p\leq\mathbf{d})^{\flat}_{\mathsf{H}}$ $=\begin{cases}p\leq\mathbf{d}^{\uparrow}&\text{if }p\geq\mathbf{d}^{\downarrow}\in{\mathsf{H}}\\ p>\mathbf{d}^{\uparrow}&\text{otherwise}\end{cases}$
$(p<\mathbf{d})^{\flat}_{\mathsf{H}}$ $=\begin{cases}p\leq\mathbf{d}&\text{if }p\leq\mathbf{d}\in{\mathsf{H}}\\ p\leq\mathbf{d}^{\uparrow}&\text{if }p\leq\mathbf{d}^{\downarrow}\in{\mathsf{H}}\\ p<\mathbf{d}^{\uparrow}&\text{otherwise}\end{cases}$
@@
$\Gamma$ $=\left\{p\vee\neg p\mid p\in\mathsf{Pr}(\chi\wedge q)\right\}\cup\{q\}$
${\mathsf{H}}$ $=\{p\geq\mathbf{1}\mid p\in\mathsf{Pr}(\chi)\}\cup\{p\leq\mathbf{0}\mid p\in\mathsf{Pr}(\chi)\}$
@@
$v^{\mathsf{cl}}(r)$ $=\begin{cases}1&\text{if }v(s\geq\mathbf{1})=1\text{ or }v(r\leq\mathbf{0})=0\\ 0&\text{if }v(r\leq\mathbf{0})=1\text{ or }v(r\geq\mathbf{1})=0\end{cases}$
@@
$\tau^{\flat}_{\lambda_{k}}$ $=\begin{cases}\bigodot\limits_{{\begin{matrix}\lambda^{\prime}\in\tau\\ \lambda^{\prime}\neq\lambda\end{matrix}}}\lambda^{\prime}\odot{\lambda^{\flat}_{k}}_{\mathsf{H}}&\text{if }{\lambda^{\flat}_{k}}_{\mathsf{H}}\text{ is defined}\\ \bigodot\limits^{k-1}_{i=1}\lambda_{i}\odot\bigodot\limits^{n}_{j=k+1}\!\!\!\!\lambda_{j}&\text{otherwise}\end{cases}$
@@ Appendix C Proofs of Section 5
$\Gamma$ $=\{\kappa_{i}\mid i\in\{1,\ldots,m\}\}\cup\{\tau_{j}\mid j\in\{1,\ldots,n\}\}$
@@
$\Gamma^{\mathbb{LI}}$ $=\{\kappa^{\mathbb{LI}}_{i}\!\mid i\!\in\!\{1,\ldots,m\}\}\!\cup\!\bigcup\limits^{n}_{j=1}\{(\lambda^{j}_{i})^{\mathbb{LI}}\!\mid i\!\in\!\{1,\ldots,k_{j}\}\}$
@@
$\Gamma_{\mathbf{int}}$ $=\left\{\bigoplus\limits_{q_{i}\in\kappa_{i}}(q_{i}\!\geq\!\mathbf{1})\oplus\bigoplus\limits_{\neg r_{i}\in\kappa_{i}}(r_{i}\!<\!\mathbf{1})\mid i\in\{1,\ldots,n\}\right\}$
${\mathsf{H}}_{\mathbf{int}}$ $=\{s\geq\mathbf{1}\mid s\in{\mathsf{H}}\}\cup\{t<\mathbf{1}\mid\neg t\in{\mathsf{H}}\}$
@@
For the membership, we construct a reduction to the classical Horn abduction for which solution recognition is polynomial. Let $\Gamma=\{\kappa_{1},\ldots,\kappa_{m},\kappa^{\prime}_{1},\ldots\kappa^{\prime}_{m^{\prime}}\}$ be a set of interval clauses of the form $\kappa_{i}=\bigodot_{j=1}^{n_{i}}(p_{j}\lozenge\mathbf{c}_{j})\rightarrow(q_{i}\lozenge\mathbf{d}_{i})$ and $\kappa^{\prime}_{i}=\bigodot_{j=1}^{n^{\prime}_{i}}(p^{\prime}_{j}\lozenge\mathbf{c}_{j})\rightarrow\bot$.
Now, for each interval literal $\lambda$, let $r_{\lambda}$ be a fresh propositional variable. We define
$\kappa_{\mathcal{H}}$ $\coloneqq\bigwedge_{j=1}^{n_{i}}r_{p_{j}\lozenge\mathbf{c}_{j}}\rightarrow r_{q_{i}\lozenge\mathbf{d}_{i}}$ $\kappa^{\prime}_{\mathcal{H}}$ $\coloneqq\bigwedge_{j=1}^{n^{\prime}_{i}}r_{p^{\prime}_{j}\lozenge\mathbf{c}_{j}}\rightarrow\bot$
We consider the following classical abduction problem: $\mathbb{P}_{\mathcal{H}}=\langle\Gamma_{\mathcal{H}},\tau_{\mathcal{H}},{\mathsf{H}}_{\mathcal{H}}\rangle$ defined as described below.
$\Gamma_{\mathcal{H}}$ $=\left\{\kappa_{\mathcal{H}}\mid\kappa\in\Gamma\right\}\cup\{\kappa^{\prime}_{\mathcal{H}}\mid\kappa^{\prime}\in\Gamma\}$
[…]
@@
${\mathsf{H}}_{\mathcal{H}}$ $=\{r_{\lambda}\mid\lambda\in{\mathsf{H}}\}$ (9)
As one can see, $\Gamma_{\mathcal{H}}$ is polynomial in the size of $\Gamma$ and (recall Proposition 3) it takes polynomial time to determine whether $\lambda\rightarrow\lambda^{\prime}$ is Ł-valid and $\lambda\odot\lambda^{\prime}$ is Ł-unsatisfieable. Furthermore, $\Gamma_{\mathcal{H}}$ is a set of Horn formulas and only contains variables of the form $r_{\lambda}$ with $\lambda$ being an interval term occurring in $\mathbb{P}$. We show that $\sigma$ is a (proper) solution to $\mathbb{P}$ iff $\sigma_{\mathcal{H}}=\bigwedge\limits_{\lambda\in\sigma}r_{\lambda}$ is a (proper) solution to $\mathbb{P}_{\mathcal{H}}$.
Assume that $\sigma\!\in\!\mathcal{S}(\mathbb{P})$. We show that $\Gamma_{\mathcal{H}},\sigma_{\mathcal{H}}\!\models^{\mathsf{cons}}_{\mathsf{CPL}}\!\tau_{\mathcal{H}}$. Let $v$ be an Ł-valuation witnessing that $\Gamma,\sigma\not\models_{{\textsf{\L }}}\bot$. We construct the classical valuation $v_{\mathcal{H}}$ as follows:
$v_{\mathcal{H}}(r_{\lambda})$ $=\begin{cases}1&\text{if }v(\lambda)=1\\ 0&\text{otherwise}\end{cases}$ (10)